40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 10 15:54:14 2015, Steve, Update, VAC, Vacuum comp. rebooted afterReb061015.png
    Reply  Thu Jun 11 19:40:59 2015, Koji, Update, VAC, Vacuum comp. rebooted 
       Reply  Fri Jun 12 08:40:17 2015, Steve, Update, VAC, Vacuum comp. rebooted vacMonNoGauges.pngprepReboot.png
Message ID: 11352     Entry time: Wed Jun 10 15:54:14 2015     Reply to this: 11353
Author: Steve 
Type: Update 
Category: VAC 
Subject: Vacuum comp. rebooted 

Koji and Steve succeded rebooting C1vac1, C1vac2 and pressure reading is working now

More tomorrow .........

 

Attachment 1: afterReb061015.png  34 kB  | Hide | Hide all
afterReb061015.png
ELOG V3.1.3-